OpenStudy (anonymous):

In the Savannah, a sinkhole opened catching some ostrich and cheetah. There was a total of 13 animals to be rescued. If 44 INDIVIDUAL feet were pulled from the muck, how many ostrich and cheetah were rescued?

OpenStudy (anonymous):

If you are going to provide me the answere to this problem PLEASE show your work lol i need the visual

OpenStudy (anonymous):

someone please help me with this

OpenStudy (anonymous):

4x + 2y = 44 x + y = 13 where x = cheetahs and y = ostrich

OpenStudy (anonymous):

solve this system of equation and u have it

OpenStudy (anonymous):

need any more help?

OpenStudy (anonymous):

but i dont know how many cheetah or ostrich there are so what do i plug in for x an y?

OpenStudy (anonymous):

u solve for x and y 4x + 2y = 44 x + y = 13 multiply 2nd equn. by 2 2x + 2y = 26 subtract this from first equn.: 2x = 18 x = 9 = number of cheetahs 9 + y = 13 y = 4 so 9 cheetahs and 4 ostriches is your answer

OpenStudy (anonymous):

ok i understand now thank you so much for your help!! : D

OpenStudy (anonymous):

at first i shld perhaps have been a bit more clear and said let number of cheetahs = x and number of ostriches = y - that made 4x cheetah feet and 2x ostrich feet and that = 44. total number of animals = x + y = 13

OpenStudy (anonymous):

no probs oh - please click good answer to close question

OpenStudy (anonymous):

lol its ok i still was able to understand , and no problem!

OpenStudy (anonymous):

fine